# TADA: Improved Diffusion Sampling with Training-free Augmented DynAmics

### v1

@@ D.1 General $N$ variable dynamics
Appendix E Detailed Explanations
E.1 Explicit form of ${\mathbf{A}}_{t}$ and ${\mathbf{b}}_{t}$
@@
AGM[9] $\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$ $[0,1]^{\mathsf{T}}$ $\frac{-4}{t-1}\left(\frac{x1-{\mathbf{x}}_{t}^{(0)}}{1-t}-{\mathbf{x}}_{t}^{(1)}\right)$
CLD[9] $\begin{bmatrix}0&-M^{-1}\\ 1&\Gamma M^{-1}\\ \end{bmatrix}\beta$ $[0,\Gamma\beta]^{\mathsf{T}}$ $\nabla_{{\mathbf{x}}_{t}^{(1)}}\log p({\mathbf{x}},t)$
E.2 Explicit form of mean and convariance matrix
@@
Eq.51 is a homogeneous Lyapunov ODE whose unique solution is exactly (see Appendix.G for more details):
@@
E.3 Previous Fast Solver
@@
Heun[17] Single Step $2$ N/A N/A $\sigma_{t}$
@@
E.4 Extended Flow Matching
@@
E.5 Degenerate Case of TADA
@@
Appendix F Additional Qualitative Comparision
@@
Appendix G Solution of the homogeneous Lyapunov ODE
@@
Appendix H Correlation between two Guassian Variable
Lemma H.1.
@@ Appendix H Correlation between two Guassian Variable
Because $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},I_{d})$ and the projector onto ${\mathbf{g}}_{t}$ is orthogonal to the projector onto the complement, ${\mathbf{g}}_{t}^{\mathsf{T}}\bm{\epsilon}$ and ϵsubscriptbold-italic-$\bm{\epsilon}_{\!\perp}$ are independent Gaussian variables.
@@
Appendix I General $N$ variable MDM loss

### v2

@@ E.1 General $N$ variable dynamics
Appendix F Detailed Explanations
F.1 Explicit form of ${\mathbf{A}}_{t}$ and ${\mathbf{b}}_{t}$
@@
AGM [9] $\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$ $0,1]^{\mathsf{T}}$ $\frac{-4}{t-1}\left(\frac{x1-{\mathbf{x}}_{t}^{(0)}}{1-t}-{\mathbf{x}}_{t}^{(1)}\right)$
CLD [9] $\begin{bmatrix}0&-M^{-1}\\ 1&\Gamma M^{-1}\\ \end{bmatrix}\beta$ $0,\Gamma\beta]^{\mathsf{T}}$ $\nabla_{{\mathbf{x}}_{t}^{(1)}}\log p({\mathbf{x}},t)$
F.2 Explicit form of mean and convariance matrix
@@
Eq. 51 is a homogeneous Lyapunov ODE whose unique solution is exactly (see Appendix. H for more details):
@@
F.3 Previous Fast Solver
@@
Heun [17] Single Step 22 N/A N/A $\sigma_{t}$
@@
F.4 Extended Flow Matching
@@
F.5 Degenerate Case of TADA
@@
Appendix G Additional Qualitative Comparision
@@
Appendix H Solution of the homogeneous Lyapunov ODE
@@
Appendix I Correlation between two Gaussian Variable
Lemma I.1.
@@ Appendix I Correlation between two Gaussian Variable
Because $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},I_{d})$ and the projector onto ${\mathbf{g}}_{t}$ is orthogonal to the projector onto the complement, ${\mathbf{g}}_{t}^{\mathsf{T}}\bm{\epsilon}$ and $\bm{\epsilon}_{\!\perp}$ are independent Gaussian variables.
@@
Appendix J General $N$ variable MDM loss